OpenStudy (anonymous):

Find the exact value of the following function. Show your work and explain the steps you take to find the value. sec(17π/6)

OpenStudy (zehanz):

First, because sec is a periodic function with period 2π, subtract 2π=12π/6 from 17π/6. So you have: sec(5π/6). Now, \(sec(x)=\dfrac{1}{\cos x}\), you have \(\dfrac{1}{\cos \frac{5}{6}\pi}\). Hopefully, you have learnt this value of the cosine function by heart, so you can give the answer now.

OpenStudy (zehanz):

If that is not the case, you have to come up with \(\cos\frac{5}{6}\pi\). You can read it off this unit circle:

OpenStudy (zehanz):

So, you have seen it is \(-\dfrac{1}{2}\sqrt{3}\). Last task: you have to calculate 1 divided by this number...